Find the polar equation for x^2+y^2=2x then show it is an equation of a circle by sketching the graph
Answer #1

40 21ST WEEK TIVNI a x² + y2 = 2x we know that, X Socoso = Busino x² + y2 = 1² and o = tam /41 so Rolar equation will be. t2

answered by: ANURANJAN SARSAM